Solve
x + y - z = 0
2x - y + z = 6
-x + y + z = 8
Show your answers as (x,y,z)

Answer by vertciel(183)

Hello there,
First of all, I recategorised your question.
Now, with the math, you have a system of three equations here. You want to find what x, y, and z are so you have to get rid of the other words you don't want to find temporarily.
x + y - z = 0 -- (1)
2x - y + z = 6 -- (2)
-x + y + z = 8 -- (3)
Let's add or multiply the equations for which we can eliminate things the easier.
(1) + (3): 2y = 8 --> y = 4
(1) + (2): 3x = 6 --> x = 2
I found x and y for you. Help yourself into finding z.